OpenStudy (anonymous):

Devaughn is 14 years older than Sydney. The sum of their ages is 94 . What is Sydney's age?

OpenStudy (johnweldon1993):

Lets let Devaughn's age be 'a' And sydney be 's' So we know together they sum to 94 a + s = 94 But we also know...that Devaughn is 14 years older than sydney...so a = s + 14 So now tht we know what 'a' equals...we can plug it into our first equation we made a + s = 94 will become s + 14 + s = 94

OpenStudy (anonymous):

40 years old! thank you

OpenStudy (rob1525):

that's right john...Amanda just solve for s then for a

OpenStudy (johnweldon1993):

We can simplify and make this 2s + 14 = 94 And then use algebra to solve for 's' **which yes is 40 lol great job!**

OpenStudy (anonymous):

thanks!!!

OpenStudy (anonymous):

these word problems are frustrating

OpenStudy (johnweldon1993):

it's all a matter of setting em up...after that like you can see, it's easy :)

OpenStudy (anonymous):

wow, i just understood the next one i did because of how you broke it down for me, thank you so much!

OpenStudy (anonymous):

actually this one has me confused, tal of 357 tickets were sold for the school play. They were either adult tickets or student tickets. The number of student tickets sold was two times the number of adult tickets sold. How many adult tickets were sold?

OpenStudy (johnweldon1993):

Well good that it helped...now for this one So we can let ADULT tickets = 'a' and STUDENT tickets = 's' The total was 357 so a + s = 357 And we know that there were 2 times as many student tickets sold than adults...so s = 2a Now we just plug it in \[\large a + 2a =357\]